OpenStudy (anonymous):

I just want to clear up something to do with factorising indices. Can someone just spend a little time for me :)

OpenStudy (anonymous):

Which term would you factor out? Because you can take out 2^2n or 2^n-1. Which one should you take out? Or its just random \[\huge\frac{2^{2n}-2^{n-1} }{ 2^{n}-2^{-1} }\]

OpenStudy (anonymous):

for the numerator it would be \[\huge2^{2n}(1 - 2^{-n-1})\] Would it?

OpenStudy (anonymous):

What if you factor out just \(2^n\)

OpenStudy (anonymous):

well ill try  this is right?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \huge\frac{2^{2n}-2^{n-1} }{ 2^{n}-2^{-1} } \] \[ \huge\frac{2^n(2^{n}-2^{-1} )}{ 2^{n}-2^{-1} } \]

OpenStudy (anonymous):

oh yes thats right :)

OpenStudy (anonymous):

so how did you know to factor out 2^n? You just looked at it?

OpenStudy (anonymous):

I wanted something similar to the bottom.

OpenStudy (anonymous):

so you keep thinking of a number to factor out till you can simplify?

OpenStudy (kuoministers):

you should be able to think a good number to factor out when you have experience :)